OpenStudy (anonymous):

help me please I will medal and fan ! What is the y-intercept of this line? y = 5x + 8

OpenStudy (kobeni-chan):

The y-intercept can be found when x=0. Substitute the value of 0 for x, then solve.

OpenStudy (anonymous):

it is acually 1.6Y

OpenStudy (kobeni-chan):

y = 5(0) + 8 y = 0 + 8 y = 8

OpenStudy (anonymous):

yea ur right @kobeni-chan i forget to check my math

OpenStudy (kobeni-chan):

lol that's okay @edwards555

OpenStudy (anonymous):

k

OpenStudy (k_lynn):

Right. When you want to find the Y intercept of an equation, you will plug in 0 for x and solve. You will do the opposite if you want to find x.